OpenStudy (anonymous):

Earth travels in an approximately circular orbit around the sun. The average radius of Earth's orbit around the sun is 9.3 x 10^7miles. Earth takes one year, or 365 days, to complete one orbit. a. Use the formula for the circumference of the circle to determine the distance,L, the Earth travels in one year. Express your answers in scientific notation. L= a x 10^n miles, where a= ___and n=___ . b. How many hours, H, are in one year? H= ___hours c. Speed is distance divided by time. Find the orbital speed, V, of Earth in miles per hour. V= ____ miles per hour

OpenStudy (anonymous):

for part b 365 days* 24hrs/day

OpenStudy (anonymous):

Oh yeah I just catched that ^.^

OpenStudy (anonymous):

not sure about the other 2 sorry part c could be 9.3 x 10^7/part b

OpenStudy (anonymous):

Thank you:) I will try solving the problem that way....

OpenStudy (mathstudent55):

Part a. C = 2(pi)r The circumference of a circle is 2 times pi times the radius of the circle.

OpenStudy (mathstudent55):

C = 2(pi)r = 2 * 3.14 * 9.3 x 10^7 = 5.8 x 10^8 L = 5.8 x 10^8 miles

OpenStudy (mathstudent55):

b. In one year there are 365 days. In one day there are 24 hours. 1 year * 365 days/year * 24 hours/day = 8,760 hours

OpenStudy (anonymous):

It work. Thank you.

OpenStudy (mathstudent55):

Did you get part c?

OpenStudy (anonymous):

No, not yet...I still keep getting it wrong:/

OpenStudy (mathstudent55):

For part c, divide the answer of part a by the answer of part b. Speed = 5.8 x 10^8 miles / 8,760 hours

OpenStudy (anonymous):

yes I figured it out after four tries....thank you!!!!!!!!!!!! I saw it because of the untis miles per hour XD

OpenStudy (mathstudent55):

great

OpenStudy (anonymous):

Aww..thank you for the medal, but I still don't understand what they are for.

OpenStudy (mathstudent55):

Me neither, but what the heck.